### v4

# LIA: Privacy-Preserving Data Quality Evaluation in Federated Learning Using a Lazy Influence Approximation

Abstract
In Federated Learning, it is crucial to handle low-quality, corrupted, or malicious data. However, traditional data valuation methods are not suitable due to privacy concerns. To address this, we propose a simple yet effective approach that utilizes a new influence approximation called ”lazy influence” to filter and score data while preserving privacy. To do this, each participant uses their own data to estimate the influence of another participant’s batch and sends a differentially private obfuscated score to the central coordinator. Our method has been shown to successfully filter out biased and corrupted data in various simulated and real-world settings, achieving a recall rate of over $>90\%$ (sometimes up to $100\%$) while maintaining strong differential privacy guarantees with $\varepsilon\leq 1$.
Index Terms:
Federated Learning, Data Filtering, Data Poisoning, Robustness, Privacy.
I Introduction
The success of Machine Learning (ML) depends to a large extent on the availability of high-quality data [49]. Having the ability to score and filter irrelevant, noisy, or malicious data can significantly improve model accuracy and speed up training. However, ensuring data quality is especially challenging in Federated Learning (FL) [45, 32, 57]. In FL, a single Center uses data from independent and sometimes self-interested data holders to train a model jointly. Data holders often operate resource-constrained edge devices and include businesses and medical institutions that must protect the privacy of their data due to confidentiality or legal constraints. Thus, we never have direct access to the training data in order to inspect it, and privacy is a strong concern. This raises our main question: how can we evaluate data quality in a distributed manner while preserving privacy?
Imagine a collaborative project involving multiple hospitals across the globe, each with its own repository of chest X-ray scans of patients. The goal is to use FL to train a robust ML model capable of diagnosing heart conditions from these X-rays [51]. However, each hospital adheres to different data collection protocols. E.g., Hospital A may use a different scanner making it incompatible with the rest of the data. If included, it can lead to a large drop in accuracy [42]. More insidiously, Hospital B may habitually include textual annotations directly on the scans for internal communication. The model may rely on these shortcuts instead of looking at the actual scan [48]. At the same time, it is important not to discard data about rare diseases, even if it might appear very different from most other data.
To address these issues, we propose a “rite of passage” mechanism where a hospital’s data is admitted into the FL system only if no issues are detected. By this, we not only ensure the integrity of the FL procedure, but can also flag potentially bad data-collection procedures for the hospitals. However, note that patient X-rays are very sensitive information and protected by strong privacy regulations such as HIPAA. Thus we need to ensure that both the model training and (for our purpose) more importantly the data quality assessment is performed in a federated and privacy preserving manner minimizing the patient information being leaked.
To prevent such a scheme from filtering out valuable data about rare conditions, quality assessment cannot be based on simple data similarity, which would consider them outliers. A clean way of quantifying the quality of data point(s) is via the notion of influence [35, 11]. Intuitively, influence quantifies the marginal contribution of a data point (or batch of points) on a model’s accuracy. One can compute this by comparing the difference in the model’s empirical risk when trained with and without the point in question. While the influence metric can be highly informative, it is impractical to compute: re-training a model is time-consuming, costly, and often impossible, as participants do not have access to the entire dataset. We propose a simple and practical approximation of the sign of the exact influence (Lazy Influence Approximation) which is based on an estimate of the direction of the model after a small number of local training epochs with the new data.
Figure 1: Data filtering procedure. Data contributors $A_{1}$ and $A_{2}$ want to join the federation, but might have biased or corrupted data (e.g., watermarks on $A_{1}$ X-rays). $A_{1}$ sends a differentially private (DP) partially updated model ($\tilde{\theta}_{t}^{A_{1}}$) to validators $B_{i}$, who submit a DP vote based on the performance they observe when using the model $\tilde{\theta}_{t}^{A_{1}}$ on their task (Lazy Influence Approximation). The aggregated votes are used as a ‘rite of passage’, to decide whether to incorporate $A_{1}$’s data. The same process happens for $A_{2}$. $A_{2}$ is accepted in the federation, while $A_{1}$ is filtered out. Filtering can significantly improve the model’s accuracy from 61% (no filtering) to 73% (proposed), matching the performance of the optimal (oracle) filtering on diagnosing heart conditions from X-rays using real-data from [51].
The second challenge is to limit the information leaked by this approximate influence measure to preserve the privacy of the data. Many approaches to Federated Learning (e.g., [46, 55]) remedy this by combining FL with Differential Privacy (DP) [17, 19, 20, 31, 15, 9], a data anonymization technique that many researchers view as the gold standard [54]. We show how the sign of influence can be approximated in an FL setting while maintaining strong differential privacy guarantees. Specifically, there are two sets of participants’ data that we need to protect: the training and the validation data (see also Section I-B). We use DP-SGD for this training to ensure datapoint-level local DP [1, 66] To ensure the privacy of the validation data and the influence approximation itself, we employ a differentially private defense mechanism based on the idea of randomized response [59] (inspired by [22]). Together the two mechanisms ensure strong, worst-case differential privacy guarantees while allowing for accurate data filtering. See Figure 1 for an overview of our approach.
I-A Our Contributions
We address two major challenges in this work: (i) efficiently estimating the quality of a batch of training data, and (ii) keeping both the training and validation data used for this estimate private. For the former, we develop a novel metric called Lazy Influence Approximation (LIA), while for the latter, we combine DP-SGD and a differentially-private voting scheme. More specifically:
(1) We present a novel technique (Lazy Influence Approximation) for scoring and filtering data in Federated Learning.
(2) We show that our proposed distributed influence aggregation scheme allows for robust scoring, even under rigorous, worst-case differential privacy guarantees (privacy cost $\varepsilon<1$). This is the recommended value in DP literature and much smaller than many other AI or ML applications. [^1] AI or ML applications often assume $\varepsilon$ as large as 10 [55] (see, e.g., [53]). For specific attacks, $\varepsilon=10$ means that an adversary can theoretically reach an accuracy of $99.99\%$ [55].
(3) We evaluate our approach on four well-established datasets: CIFAR10, CIFAR100, real-data on Human Activity Recognition on edge devices [3], and real-data on diagnosing heart conditions from X-rays [51]. Our simulations include two different modalities (image and sensor data), and corruption of both the input $X$, and the labels $y$. We demonstrate that filtering using our scheme can eliminate the adverse effects of inaccurate data. We also show it can be easily combined with other robust-FL training methods, to further enhance the model’s performance.
I-B High-Level Description of Our Setting
A center $C$ coordinates a set of participants that contribute their data to train a single model (Figure 1). $C$ has a small set of ‘warm-up’ data, which are used to train an initial model $M_{0}$ that captures the desired input/output relation. The model is then updated with the contributions of the participants. We assume that each data holder participant has a set of training points that will be used to improve the model and a set of validation points that will be used to evaluate other participants’ contributions. It must be kept private to prohibit participants from tailoring their contributions to the validation data. Each data holder participant can assume two roles: the role of the contributor ($A$) and the role of the validator ($B$). As a contributor, a participant performs a small number of local epochs to $M_{t}$ – enough to get an estimate of the gradient [^2] The number of local epochs is a hyperparameter. We do not need to train the model fully (e.g., 2-5 epochs). – using a batch of his training data $z_{A,t}$. Subsequently, $A$ sends the updated partial model $M_{t,A}$, with specifically crafted noise to ensure DP, to the validators. The applied noise protects the updated gradient while still retaining information on the usefulness of data. Each validator $B$ uses its validation dataset to approximate the empirical risk of $A$’s training batch (i.e., the approximate influence). This is done by evaluating each validation point and comparing the loss. In an FL setting, we can not re-train the model to compute the exact influence; instead, $B$ performs only a small number of training epochs, enough to estimate the direction of the model (Lazy Influence Approximation). As such, we look at the sign of the approximate influence (and not the magnitude). Each validator aggregates the signs of the influence for each validation point, applies controlled noise to ensure DP, and sends this information to the center. Finally, the center accepts $A$’s training batch if most of the $B$s report positive influence and rejects otherwise (majority voting). To recognize the value of data about rare conditions, it is important to use a broad sample of validation data. Following the federated learning idea, we distribute evaluation among many participants. Anyone capable of using the model can be a validator, so there could be many more than contributors.
II Related Work and Discussion
Federated Learning
Federated Learning (FL) [45, 32, 57, 38] has emerged as an alternative method to train ML models on data obtained by many different agents. In FL, a center coordinates agents who acquire data and provide model updates. FL has been receiving increasing attention in both academia [39, 64, 26, 6] and industry [25, 8], with a plethora of real-world applications.
Influence functions
Influence functions are a standard method from robust statistics [11] (see also Section III), which were recently used as a method of explaining the predictions of black-box models [35]. They have also been used in the context of fast cross-validation in kernel methods and model robustness [43, 10]. While a powerful tool, computing the influence involves too much computation and communication, and it requires access to the training and validation data (see [35] and Section III). There has also been recent work trying to combine Federated Learning with influence functions [62], though to the best of our knowledge, we are the first to provide a privacy-preserving alternative.
Data Filtering
A common but computationally expensive approach for filtering in ML is to use the Shapley Value of the Influence to evaluate the quality of data [30, 23, 29, 63, 24, 60]. Other work includes, for example, rule-based filtering of least influential points [50], or constructing weighted data subsets (corsets) [14]. Because of the privacy requirements in FL, contributed data is not directly accessible for assessing its quality. [56] propose a decentralized filtering process specific to federated learning, yet they do not provide any formal privacy guarantees. While data filtering might not always pose a significant problem in traditional ML, in an FL setting, it is more important because even a small percentage of mislabeled data can result in a significant drop in the combined model’s accuracy.
Client Selection and Attack Detection
Our setting can also be interpreted as potentially adversarial, but it should not be confused with Byzantine robustness. We do not consider threat scenarios as described in [7] and [52], where participants carefully craft malicious updates. Instead, we assume that the data used for those updates might be corrupt. For completeness and in lack of more relevant baselines, we compare our work to two Byzantine robust methods: KRUM [5], Trimmed-mean [65], and Centered-Clipping [33] (along to an oracle filter). These methods, though, require gradients to be transmitted as is, i.e., they lack any formal privacy guarantee to the participants’ training data. Furthermore, both of these techniques require the center to know the number of malicious participants a priori. Nevertheless, we show that our approach can be combined with such robust-FL training methods, to further enhance the model’s performance.
Differential Privacy
Differential Privacy (DP) [17, 19, 20] has emerged as the de facto standard for protecting the privacy of individuals. Informally, DP captures the increased risk to an individual’s privacy incurred by participating in the learning process. Consider a participant being surveyed on a sensitive topic as a simplified, intuitive example. To achieve differential privacy, one needs a source of randomness; thus, the participant decides to flip a coin. Depending on the result (heads or tails), the participant can reply truthfully or randomly. An attacker can not know if the decision was taken based on the participant’s preference or due to the coin toss. Of course, to get meaningful results, we need to bias the coin toward the actual data. In this simple example, the logarithm of the ratio $Pr[\text{heads}]/Pr[\text{tails}]$ represents the privacy cost (also referred to as the privacy budget), denoted traditionally by $\varepsilon$. Yet, one must be careful in designing a DP mechanism, as it is often hard to practically achieve a meaningful privacy guarantee (i.e., avoid adding a lot of noise and maintain high accuracy) [55, 13]. A variation of DP, instrumental in our context, given the decentralized nature of federated learning, is Local Differential Privacy (LDP) [34, 21]. LDP is a generalization of DP that assumes all communication between the data providers and the Center is public. Hence, we do not rely on the Center to take appropriate privacy measures and instead try limit information leakage in all communication. This allows us to work with untrusted central coordinators. An orthogonal concept is datapoint-level vs. user-level DP [58, 4, 44]. In user-level privacy, we attempt to protect datasets at the level of the data provider, rather than at the level of data-point. While this is typical in the cross-device FL where each data-provider represents the data of a single user, this does not make sense for multi-institutional collaboration [41]. In the setting we consider, each hospital (data provider) aggregates data from many patients. Thus, each user we want to protect is associated with a single data point, making datapoint-level DP more relevant. Thus, we consider datapoint-level local DP. We assume that the participants and the Center are honest but curious, i.e., they don’t actively attempt to corrupt the protocol but will try to learn about each other’s data.
III Methodology
We aim to address two challenges: (i) approximating the influence of a (batch of) data point(s) without having to re-train the entire model from scratch and (ii) doing so while protecting the privacy of the training and validation data. The latter is essential not only to protect users’ sensitive information but also to ensure that malicious participants can not tailor their contributions to the validation data. We first introduce the notion of influence [11] and our proposed lazy approximation. Second, we describe a differentially private voting scheme for crowdsourcing the approximate influence values.
Setting
We consider a classification problem from some input space $\mathcal{X}$ (e.g., features, images, etc.) to an output space $\mathcal{Y}$ (e.g., labels). In a FL setting, there is a center $C$ that wants to learn a model $M(\theta)$ parameterized by $\theta\in\Theta$, with a non-negative loss function $L(z,\theta)$ on a sample $z=(\bar{x},y)\in\mathcal{X}\times\mathcal{Y}$. Let $R(Z,\theta)=\frac{1}{n}\sum_{i=1}^{n}L(z_{i},\theta)$ denote the empirical risk, given a set of data $Z=\{z_{i}\}_{i=1}^{n}$. We assume that the empirical risk is differentiable in $\theta$.
Definitions
In simple terms, influence measures the marginal contribution of a data point on a model’s accuracy. A positive influence value indicates that a data point improves model accuracy, and vice-versa. More specifically, let $Z=\{z_{i}\}_{i=1}^{n}$, $Z_{+j}=Z\cup z_{j}$ where $z_{j}\not\in Z$, and let $\hat{R}=\min_{\theta}R(Z,\theta)$ and $\hat{R}_{+j}=\min_{\theta}R(Z_{+j},\theta)$, where $\hat{R}$ and $\hat{R}_{+j}$ denote the minimum empirical risk of their respective set of data. The influence of datapoint $z_{j}$ on $Z$ is defined as $\mathcal{I}(z_{j},Z)\triangleq\hat{R}-\hat{R}_{+j}$.
III-A Shortcomings of the Exact and Approximate Influence in a FL Setting
Despite being highly informative, influence functions have not achieved widespread use in FL (or ML in general). This is mainly due to the computational cost. The exact influence requires complete retraining of the model, which is time-consuming and very costly, especially for state-of-the-art, large ML models (importantly for our setting, we do not have direct access to the training data). Recently, the first-order Taylor approximation of influence [35] (based on [12]) has been proposed as a practical method to understanding the effects of training points on the predictions of a centralized ML model. While it can be computed without having to re-train the model, according to the following equation $\mathcal{I}_{appr}(z_{j},z_{val})\triangleq-\nabla_{\theta}L(z_{val},\hat{\theta})H^{-1}_{\hat{\theta}}\nabla_{\theta}L(z_{j},\hat{\theta})$, it is still ill-matched for FL models for several key reasons.
Computing the influence approximation of [35] requires forming and inverting the Hessian of the empirical risk. With $n$ training points and $\theta\in\mathbb{R}^{m}$, this requires $O(nm^{2}+m^{3})$ operations [35], which is impractical for modern-day deep neural networks with millions of parameters. To overcome these challenges, [35] used implicit Hessian-vector products (HVPs) to more efficiently approximate $\nabla_{\theta}L(z_{val},\hat{\theta})H^{-1}_{\hat{\theta}}$, which is typically $O(m)$. While this is a somewhat more efficient computation, it is communication-intensive, as it requires transferring all of the (either training or validation) data at each FL round. Most importantly, it can not provide any privacy to the users’ data, an important, inherent requirement/constraint in FL. Finally, the loss function has to be strictly convex and twice differentiable (which is not always the case in modern ML applications). [35] proposed to swap out non-differentiable components for smoothed approximations, but there is no quality guarantee of the influence calculated in this way.
III-B Lazy Influence (LIA): A Practical Influence Metric for Filtering Data in FL Applications
The key idea is that we do not need to approximate the influence value to filter data; we only need an accurate estimate of its sign (in expectation). Recall that a positive influence value indicates a data point improves model accuracy. Thus, we only need to approximate the sign of the loss and use that information to filter out data whose influence falls below a certain threshold.
Recall that each data holder participant may assume two roles: the role of the contributor ($A$) and the role of the validator ($B$). The proposed approach works as a ‘rite of passage’ for contributor $A$, arriving at round $t$. Our approach works as follows (Algorithm 1):
(i) At federated learning round $t$ (model $M_{t}(\theta_{t})$), the contributor participant $A$ performs a small number $k$ of local epochs to $M_{t}$ using a batch of his training data $Z_{A,t}$, resulting in $\tilde{\theta}_{t}^{A}$. $k$ is a hyperparameter. $\tilde{\theta}_{t}^{A}$ is the partially trained model of participant $A$, where most of the layers, except the last one, have been frozen. The model should not be fully trained for two key reasons: efficiency and avoiding over-fitting (e.g., in our simulations, we only performed 1-20 epochs). Furthermore, $A$ adds noise to $\tilde{\theta}_{t}^{A}$ (see Section III-B2) to ensure strong, worst-case local differential privacy. Finally, $A$ sends only the last layer (to reduce communication cost) of $\tilde{\theta}_{t}^{A}$ to every other participant.
(ii) Each validator $B$ uses his validation dataset $Z_{val}^{B}$ to estimate the sign of the influence using Equation 1. Next, the validator applies noise to $\mathcal{I}_{LIA}(Z_{val}^{B})$, as described in Section III-B3, to ensure strong, worst-case differential privacy guarantees (i.e., keep his validation dataset private).
$\mathcal{I}_{LIA}(Z_{val}^{B})\triangleq\operatorname{sign}\left(\sum_{z_{val}\in Z_{val}^{B}}L(z_{val},\theta_{t})-L(z_{val},\theta_{t}^{A})\right)$ (1)
(iii) Finally, the center $C$ aggregates the obfuscated votes $\mathcal{I}_{LIA}(Z_{val}^{B})$ from all validators and filters out data with cumulative score below a threshold ($\sum_{\forall B}\mathcal{I}_{LIA}(Z_{val}^{B})<T$). Specifically, we cluster the votes into two clusters (using k-means) and use the arithmetic mean of the cluster centers as the filtration threshold.
1
2for Data contributor $A$ arriving at round $t$ do
3
4      $A$: Performs $k$ local epochs with $Z_{A,t}$ on the partially-frozen model $\tilde{\theta}_{t}^{A}$.
5      $A$: Applies DP noise to $\tilde{\theta}_{t}^{A}$.
6      $A$: sends last layer of $\tilde{\theta}_{t}^{A}$ to a subset of the participants, that will be acting as validators.
7      for $B_{i}$ in $Validators(t)$ do
8             $B_{i}$: Evaluates the loss of $Z_{val}^{B_{i}}$ on $\theta_{t}$ (current joint model of the federation)
9            $B_{i}$: Evaluates the loss of $Z_{val}^{B}$ on $\tilde{\theta}_{t}^{A}$ (partially trained model from $A$)
10            $B_{i}$: Calculates vote $v$ (Lazy Influence Approximation), according to Equation 1
11            $B_{i}$: Applies noise to $v$ according to his privacy parameter $p$ to get $v^{\prime}$ (Equation 2)
12            $B_{i}$: Sends $v^{\prime}$ to $C$ (private majority voting)
13
14      $C$: Filters out (‘rite of passage’) $A$’s data if $\sum_{\forall B_{i}}\mathcal{I}_{LIA}(Z_{val}^{B_{i}})<T$.
Algorithm 1 Filtering Poor Data Using the Lazy Influence Approximation (LIA) in FL
III-B1 Advantages of the proposed lazy influence
Depending on the application, the designer may select any optimizer to perform the model updates. We do not require the loss function to be twice differentiable and convex; only once differentiable. It is significantly more computation and communication efficient; an essential prerequisite for any FL application. This is because participant $A$ only needs to send (a small part of) the model parameters $\theta$, not his training data. Moreover, computing a few model updates (using, e.g., SGD or any other optimizer) is significantly faster than computing either the exact influence or an approximation due to the numerous challenges. As a concrete example, computing the proposed LIA is  30 times faster than the exact influence (on our simulations on real-data for Human Activity Recognition). Finally, and importantly, we ensure the privacy of both the train and validation dataset of every participant.
III-B2 Sharing the Partially Updated Model: Privacy, Communication Cost, and Scalability
Each contributor $A$ shares a partially trained model $\tilde{\theta}_{t}^{A}$ (see step (i) of Section III-B). It is important to stress that A only sends the last layer of the model. This has two significant benefits. First, it minimizes the impact of the differential privacy noise. We follow  [1, 66] to ensure strong datapoint-level local differential privacy guarantees by (i) imposing a bound on the gradient (using a clipping threshold $\Delta$), and (ii) adding carefully crafted Gaussian noise (parameterized by $\sigma$). Second, it reduces the communication overhead (e.g., in our CIFAR simulations, we only send $0.009\%$ of the model’s weights). Moreover, as explained in the Introduction, this communication cost will be incurred only once, when we use our approach as a ‘right of passage’ every time a participant joins the federation. Importantly, in practice one can randomly select a (sufficiently large [^3] With enough votes, the aggregated LIA signs will match the sign of the true influence value in expectation. The size of the validators set is problem specific. See Sections IV, IV-C, and Figure 4.) subset of validators, keeping the communication cost constant as we scale.
Figure 2: Visualization of the private voting scheme. The $x$-axis represents a contributor participant $A$. The $y$-axis shows the sum of all votes from all the validators, i.e., $\sum_{\forall B}\mathcal{I}_{LIA}(Z_{val}^{B})$. Figure 2(a) corresponds to the sum of true votes (no privacy) for the valiation data of each contributor on the $x$-axis, while Figure 2(b) depicts the sum of differentially private votes ($\varepsilon=1$), according to randomized reporting algorithm. Finally, Figure 2(c) shows the filtration threshold, corresponding to the arithmetic mean of the two cluster centers (computed using k-means).
III-B3 Differentially Private Reporting of the Influence
Along with the training data, we also need to ensure the privacy of the validation data used to calculate the influence. Protecting the validation data in an FL setting is critical since (i) it is an important constraint of the FL setting, (ii) participants want to keep their sensitive information (and potential means of income, e.g., in a crowdsourcing application) private, and (iii) the center wants to ensure that malicious participants can not tailor their contributions to the validation sets.
Each validator $B_{i}$ is assigned a random ID number every time they act in that capacity. Moreover, each can split their test data into multiple disjoint sets, and then use a each set only once. [^4] E.g., in our applications, each of these sets required at most 50 points for the vote to be accurate. We obfuscate their influence reports using RAPPOR [22], which results in an $\varepsilon$-differential privacy guarantee [20]. [^5] Note that RAPPOR provides significantly stronger privacy guarantees ($\delta=0$) that what is normally deployed in modern systems. Thus, in practice, the privacy loss $\varepsilon$ will be even lower if we allow $\delta>0$. The obfuscation (permanent randomized response [59]) takes as input the participant’s true influence value $v$ (binary) and privacy parameter $p$, and creates an obfuscated (noisy) reporting value $v^{\prime}$, according to Eq. 2. $p$ is a user-tunable parameter that allows the participants themselves to choose their desired level of privacy, while maintaining reliable filtering. The worst-case privacy guarantee can be computed by each participant a priori, using Eq. 3 [22].
$v^{\prime}=\begin{cases}+1,&\text{with probability $\frac{1}{2}p$}\\ -1,&\text{with probability $\frac{1}{2}p$}\\ v,&\text{with probability $1-p$}\end{cases}$ (2)
$\varepsilon=2\ln\left(\frac{1-\frac{1}{2}p}{\frac{1}{2}p}\right)$ (3)
It is important to note that in a Federated Learning application, the center $C$ aggregates the influence sign from a large number of participants. This means that even under really strict privacy guarantees, the aggregated influence signs (which is exactly what we use for filtering) will match the sign of the true value in expectation. This results in high-quality filtering, as we will demonstrate in Section IV.
We visualize the obfuscation process in Figure 2, which shows the sum of true votes ($y$-axis) for the validation data of each contributor ($x$-axis). Here we can see a clear distinction in votes between corrupted and correct batches. Most of the corrupted batches (corrupted contributor participants) take negative values, meaning that the majority of the validators voted against them. In contrast, the correct batches are close to the upper bound. Figure 2(b) demonstrates the effect of applying DP noise ($\varepsilon=1$) to the votes: differentiating between the two groups becomes more challenging. To find an effective decision threshold, we use k-means to cluster the votes into two clusters and use the arithmetic mean of the cluster centers as the filtration threshold (Figure 2(c)).
IV Evaluation Results
Dataset
We evaluated the proposed approach on four well-established datasets: CIFAR10 [36], CIFAR100 [36], real-data on Human Activity Recognition (HAR) on edge devices [3, 2], and real-data on diagnosing heart conditions from X-rays [51]. Between these datasets, we evaluate on two different modalities (image data and sensor data), and corruption of both the input X, and the labels y. Due to lack of space, we focus on CIFAR in this section. Similar results were achieved for the other datasets. Please see the supplement (XI) for detailed results.
Corruption Methods
(i) Random label: a random label is sampled for every training point. Used for the IID setting (as it does not make sense to assign a random label to a highly skewed Non-IID setting). (ii) Label shift: Every correct label is mapped to a different label and this new mapping is applied to the whole training dataset. Used in both IID and non-IID settings.
Setup
Our evaluation involves a single round of Federated Learning. A small portion of every dataset (around 1%) is selected as the ‘warm-up’ data used by the center $C$ to train the initial model $M_{0}$. Each participant has two datasets: a training batch ($Z_{A}$, see Section III-B, step (i)), which the participant uses to update the model when acting as the contributor participant, and a validation dataset ($Z_{val}^{B}$, see Section III-B, step (ii)), which the participant uses to estimate the sign of the influence when acting as a validator participant. As a concrete example, for the CIFAR simulations the ratio of these datasets is $2:1$. The training batch size is 100 (i.e., the training dataset includes 100 points, and the validation dataset consists of 50 points). This means that, e.g., for a simulation with 100 participants, each training batch is evaluated on $50\times(100-1)$ validation points, and that for each training batch (contributor participant $A$), the center collected ($100-1$) estimates of the influence sign (Equation 1). See the supplement (XI) for details on the HAR dataset. We report results when having 30% corrupt participants. Additional results (0%, 10%, 20%, 40%) can be found in the supplement XI. For each corrupted batch, we corrupted 100% of the data points (similar results were achieved with 90%). Each simulation was run 8 times. We report average values and standard deviations. The proposed approach is model-agnostic and can be used with any gradient-descent-based ML method.
Non-IID Setting
The main hurdle for FL is that not all data is IID. Heterogeneous data distributions are all but uncommon in the real world. To simulate non-IID data, we used the Dirichlet distribution to split the training dataset as in related literature [28, 40, 27, 68]. This distribution is parameterized by $\alpha$, which controls the concentration of different classes (see the supplement VII-A for a visualization). We report results for $\alpha\rightarrow 0.1$, as in related literature (e.g., [68]), see the supplement XI for the rest. This translates to a highly non-IID distribution, e.g., with just 3 classes per participant for the HAR dataset.
Baselines
We compare against four baselines. (i) Corrupted model: no sanitization (no filtering). (ii) Oracle filtration: ideal scenario where we know which participants contribute bad data. (iii) KRUM [5]: byzantine robustness technique that selects the best gradient out of the update based on a pair-wise distance metric. (iv) Trimmed-mean [65]: another byzantine robustness technique that takes the average of gradients instead of just selecting one, also based on a pair-wise distance metric. (v) Centered-Clipping [33]: state-of-the-art technique for byzantine robust aggregators. Importantly, we show that our filtration technique is not mutually exclusive with these aggregators, instead it is highly beneficial in conjunction to them (see Figure 3).
IV-A Model Performance
The proposed approach achieves high model performance, close to the model performance of the perfect (oracle) filtering ($13.6\%$ worse in the non-IID setting, and $0.1\%$ in the IID setting). Focusing on the non-IID setting (Figure 3(a)), which is the more challenging and relevant for FL, our approach achieves a $10.8\%$, $20.3\%$, and $57.5\%$ improvement over the Centered-Clipping, KRUM, and Trimmed-mean baselines, respectively, after 25 communication rounds. Interestingly, combining our approach with Centered-Clipping provides results almost as good as the Oracle filter (only $3\%$ worse), which comes to show that LIA can complement Byzantine robustness techniques. Finally, while in the IID setting all methods perform similarly (Figure 3(b)), recall that the baselines do not provide privacy guarantees.
(a) Label shift Non-IID ($\alpha=0.01$)
(b) Random Label IID
Figure 3: Model performance (model accuracy) over 25 communication rounds. 30% mislabel rate on CIFAR-10. The proposed (LIA) and oracle filters are used only once at the start (‘right of passage’ scenario). We compare a centralized model with no filtering (blue) to an FL model under perfect (oracle) filtering (orange), KRUM (red), Trimmed-mean (purple), Centered-Clipping(brown), our approach with FedAvg (green), and our approach with Centered-Clipping (pink). Note that the jagged line for KRUM is because only a single gradient is selected instead of performing FedAvg.
(a) Recall
(b) Precision
Figure 4: Recall and Precision on CIFAR 10, highly non-IID ($\alpha\rightarrow 0.1$), for increasing problem size (# of participants), and varying privacy guarantees (lower $\varepsilon$ provides stronger privacy). $\delta=10^{-5}$.
IV-B Filtration Metrics: Recall, Precision, Accuracy
Recall is the most informative metric to evaluate the efficiency of our filtering approach. Recall refers to the ratio of detected mislabeled batches over all of the mislabeled batches. Including a mislabeled batch can harm a model’s performance significantly more compared to removing an unaltered batch. Thus, achieving high recall is of paramount importance. Meanwhile, precision represents the ratio of correctly identified mislabeled batches over all batches identified as mislabeled. An additional benefit of using the proposed lazy influence metric for scoring data is that it also allows us to identify correctly labeled data, which nevertheless do not provide a significant contribution to the model. Finally, filtration accuracy refers to the ratio of correctly identified batches over all batches.
The proposed approach achieves both high recall and precision (Figure 4), despite the high degree of non-IID (low concentration of classes per participant). Notably, the metrics improve significantly as we increase the number of participants (horizontal axis, Figure 4). In simple terms, more validators mean more samples of the different distributions. Thus, ‘honest’ participants get over the filtering threshold, even in highly non-IID settings. Recall reaches $100\%$ and precision $96.48\%$ on CIFAR10 (Figure 4) by increasing the number of participants to just 500 (highly non-IID setting and under really strict worst-case privacy guarantees). Results for the IID setting are significantly better, with almost perfect recall, precision, and accuracy, even with only 100 participants.
IV-C Privacy
As expected, there is a trade-off between privacy and filtration quality (see Figure 4, vertical axis, where $\varepsilon$ refers to the privacy guarantee for both the training and validation data/participant votes). Nevertheless, Figure 4 demonstrates that our approach can provide reliable filtration, even under really strict, worst-case privacy requirements ($\varepsilon=1$, which is the recommended value in the DP literature [54], $\delta=10^{-5}$). Importantly, our decentralized framework allows each participant to compute and tune his own worst-case privacy guarantee a priori (see Section III-B3).
The privacy trade-off can be mitigated, and the quality of the filtration can be significantly improved by increasing the number of validators (Figure 4, horizontal axis). The higher the number of validators, the better the filtration (given a fixed number of corrupted participants). This is because as the number of validators increases, the aggregated influence signs (precisely what we use for filtering) will match the sign of the true influence value (in expectation). For $500$ validators, we achieve high-quality filtration even for $\varepsilon=0.75$. This is important given that in most real-world FL applications, we expect a large number of validators.
IV-D Robustness
We evaluated our approach for varying mislabeling percentage (0%, 10%, 20%, 30%, 40%), varying levels of non-IID-ness ($\alpha=0.1,1,10$), and a setting where not all data in a mislabeled batch are corrupt (90% corrupt). In all cases, the proposed approach enabled effective filtering (recall of $>90\%$). Please see the supplement VII-A for detailed results.
IV-E Minority Distributions
An important concern when filtering data is to ensure that clients with minority distributions will not be filtered out. Our results thus far demonstrate that our proposed approach achieves high quality filtering (outperforming established baselines), even on very hard non-IID settings (as low as just 3 classes per participant). Yet, these are aggregate results. To investigate specifically the case where clients have data from minority distributions, we consider the most extreme scenario, where only one client has data from a specific class (consider e.g., a hospital that treats a rare disease).
Specifically, we consider a scenario where every validator has data from all but one classes (leave-one-out, i.e., $\mathcal{Y}\setminus\{y_{out}\}$), drawn from a Dirichlet distribution, in a highly non-IID setting (as in our other simulations). Let a new incoming client $A$ have data from all class (including $y_{out}$). Let 5%, 10%, 15%, 20% of client $A$’s data belong to his unique class $y_{out}$, while the rest are drawn from classes $\mathcal{Y}\setminus\{y_{out}\}$ using the Dirichlet distribution. Figure 5 shows that, with high probability ($0.63$ for 5%, $0.54$ for 20%), client $A$ will not be filtered out. Our intuition is that client $A$ also has data that (partially) overlap with the validators (even in highly non-IID settings), and this overlap results in positive influence, in spite of their minority class data.
Figure 5: Probability that a client with data from a unique distribution (see Section IV-E) will be accepted in the federation. With high probability ($>50\%$) the client will not be filtered out, even in the extreme case of having 20% of data from a unique distribution. Results averaged over 512 runs.
V Conclusion
The quality of a model obtained through Federated Learning can only be as good as the quality of the contributed data. Mislabeled, corrupted, or even malicious data can result in a strong degradation of the performance of the model (Figure 1). However, it is challenging to ensure this quality while preserving the privacy guarantees that are a hallmark of federated learning, and at the same time ensure that data from minority classes is not discarded. We propose a practical influence approximation (‘lazy influence’) to be used as a ‘rite of passage’, allowing for effective filtering (recall of $>90\%$, and even up to 100%), while providing strict, worst-case $\varepsilon$-differential privacy guarantees ($\varepsilon\leq 1$) for both the training and validation data.